# Scaling Manipulation Learning with Visual Kinematic Chain Prediction

### v1

@@ 4.3 Ablation Studies
Specialized Agent BCT 1.1 59 31.7 12.7 5.2 1.4
VKT (Ours) 1.173 60.5 32.5 14.7 7.1 2.5
General Agent BCT 0.334 26.8 5.3 1.2 0.1 0
VKT (Ours) 1.083 58.3 29 13.5 5.6 1.9
Predicting the Full Chain versus Predicting Only the End-Effector. We compare the effects of forecasting the full chain and only the end-effector in Table 5. Table 5 shows that forecasting the full chain improves the performance on both Calvin and RLBench. However, despite the inability to capture rotations, end-effector VKT also delivers stronger performance as a general agent than BCT. This indicates that a key to building general agents across diverse environments is to predict robot movements directly in the image plane. Furthermore, the end-effector predictions can be useful in special cases when the robot’s body is largely invisible due to occlusions, or when robot models are unavailable. This suggests a broader applicability for visual kinematics forecasting.
Multi-View versus Single-View Predictions. We study the impacts of multi-view forecasting in Table 5. We remove the multi-view attention layers to create the single-view VKT. We average the low-level robot actions predicted from each view as the final prediction. Table 5 shows that VKT improves its 3D understanding significantly with our multi-view dual attention block by predicting consistent kinematic chains in different camera viewpoints.
5 Discussion and Conclusion
We have shown that visual kinematics forecasting of quasi-static robot movements can be a building block technique for developing general agents across diverse robot learning environments. However, there are still some open questions that are worth discussing. For example, we did not utilize wrist cameras in VKT because it does not take images of the robot. This can be resolved with RGBD wrist camera and virtual view projection [18]. Also, some public real-robot datasets lack robot URDF files and reliable camera parameters, which creates several challenges for applying our method. Finally, it is worth noting that one can re-purpose VKT to forecast not only robot movements, but also human poses and movements in order to leverage abundant human demonstration videos for robot learning.

### v3

@@ A.2 Additional Tables and Figures
Specialized MCIL [49] 0.41 34.4 5.8 1.1 0.2 0
GCBC [41] 1.11 (0.3) 64.7 28.4 12.2 4.9 1.3
BCT 1.36 (0.06) 63.2 39.6 18.6 9.7 4.9
VKT (Ours) 1.58 (0.06) 67.6 44.3 24.1 14.5 7.5
General BCT 0.48 (0.1) 36.4 8.6 2.6 0.4 0
VKT (Ours) 1.46 (0.08) 64.9 41.4 22.3 12.8 6.4
Figure A5: Failure Case Analysis. The predicted kinematic chain of the next step is colored in red. The expected action is shown in a green arrow. The first row shows tasks: drag with stick and stack blocks. The kinematic chain is visually correct to represent the grabbing and stacking action. However, the predicted action sometimes lacks the necessary precision. The second row shows tasks: push blocks and close microwave, which requires spatial reasoning. However, the predicted chain represents the wrong actions. The third row shows the cube transfer task. The next action is to move two arms closer with the right arm grasping the cube. The predicted chain is visually reasonable. However, the gripper state is ambiguous when the right arm faces the camera from the side. When the right arm drops the cube, the task fails. The fourth row shows the tasks: close jar and turn tap, which require rotation actions with only small changes in the visual kinematic chain. The predicted chain does not show rotation actions.